OpenStudy (anonymous):

How do you add logs with different bases? e.g. log (base 2) y + log (base 4) y = 6 Thanks in advance!

OpenStudy (saifoo.khan):

\[\frac{\log (y) }{\log(2)} + \frac{\log(y)}{\log4} = 6\]

OpenStudy (anonymous):

:D Thanks.

OpenStudy (saifoo.khan):

i dont know how to go from here. lol

OpenStudy (anonymous):

To be honest, I don't either. :P

OpenStudy (saifoo.khan):

lol, ok.. let me fig this out.

OpenStudy (anonymous):

:D That's nice of you, thanks!

OpenStudy (saifoo.khan):

GOt it! just a sec.

OpenStudy (mimi_x3):

is it: \[\large \log_2y + \log_4y = 6\] ?

OpenStudy (saifoo.khan):

yes.

OpenStudy (saifoo.khan):

\[\frac{2\log y + logy}{\log 2^2} = 6\]

OpenStudy (saifoo.khan):

where? @Mimi_x3

OpenStudy (anonymous):

According to this, the answer is supposed to be 16. D:

OpenStudy (saifoo.khan):

\[\frac{3 \log y }{2\log 2} = 6\] \[\frac32 \log(y-2) = 6\]

OpenStudy (saifoo.khan):

can u please guide what we will do after this? @Gab02282 . the answer is 16.

OpenStudy (anonymous):

I'm not sure, either. The question book just gave me the answer, but not the process. I'm sorry I can't be of any help. D:

OpenStudy (saifoo.khan):

no worries... @Mimi_x3 im correct. just can't find steps to move forward.. :P Do u know?

OpenStudy (anonymous):

Make "6" an exponent? Since they have similar logs, can it be like, log (base 10) 10^6? I don't know what I'm doing lol.

OpenStudy (anonymous):

After that, eliminate the logs and solve for y. Right?

OpenStudy (saifoo.khan):

we have to solve for y.. yes..

OpenStudy (anonymous):

//captain obvious lol, I'm honestly unsure.

OpenStudy (anonymous):

\[3\lg_{4}y/2\lg 2_{4} = 6\] \[2\lg _{4}2 = 2*1/2\] = 1 => \[3\lg _{4} y=6\] =>\[\lg _{4}y = 2\] \[y = 4^{2} = 16\]

OpenStudy (saifoo.khan):

@chemwile , mind showing the first two steps using draw tool?

OpenStudy (anonymous):

Thank you very much, @chemwile!

OpenStudy (anonymous):

vw